# ProG: A Graph Prompt Learning Benchmark

### v1

@@ 3 Benchmark Methods
In general, graph prompt learning aims to transfer pre-trained knowledge with the expectation of achieving better positive transfer effects in downstream tasks compared to classical ‘Pre-train & Fine-tune’ methods. To conduct a comprehensive comparison, we first consider using a supervised method as the baseline for judging positive transfer (For example, if a ‘pre-training and fine-tuning’ approach can not beat supervised results, one negative transfer case is observed). We introduce various pre-training methods to generate pre-trained knowledge, upon which we perform fine-tuning as the comparison for graph prompts. Finally, we apply current popular graph prompt learning methods with various pre-training methods to investigate their knowledge transferability. The details of these baselines are as follows (Further introduction can be referred to Appendix):
  • Supervised Method: In this paper, we utilize Graph Convolutional Network (GCN) [9] as the baseline, which is one of the classical and effective graph models based on graph convolutional operations [25, 37] and also the backbone for both ‘Pre-train & Fine-tune’ and graph prompt learning methods. We also explore more kinds of graph models like GraphSAGE [6], GAT [27], and Graph Transformer [20], and put the results in Appendix.
@@ 4.1 Datasets
MUTAG 188 17.9 39.6 7 2 G small molecule
IMDB-BINARY 1000 19.8 193.1 0 2 G social network
COLLAB 5000 74.5 4,914.4 0 3 G social network
@@
ENZYMES 600 32.6 124.3 3 6 G proteins
DD 1,179 284.1 14,322.4 89 2 G proteins
@@
To figure out how adaptive existing graph prompts on different graphs, we conduct our experiments across 15 datasets on node-level or graph-level tasks, providing a broad and rigorous testing ground for algorithmic comparisons. As shown in Table 4.1, we include 7 node classification datasets, covering homophilic datasets (Cora, Citeseer, PubMed, ogbn-arxiv) [19, 15, 8], heterophilic datasets (Wisconsin, Texas, Actor) [16, 26], and large-scale datasets (ogbn-arxiv). Additionally, we consider 8 graph classification datasets from different domains, including social networks (IMDB-B, COLLAB) [33], biological datasets (ENZYMES, PROTEINS, DD) [3, 1, 30], and small molecule datasets (MUTAG, COX2, BZR) [10, 18]. More details on these datasets can be found in Appendix.
@@ 4.2 Implementation Details
Metrics. Following existing multi-class classification benchmarks, we select Accuracy, Macro F1 Score, and AUROC (Area Under the Receiver Operating Characteristic Curve) as our main performance metrics. The Accuracy results are used to provide a straightforward indication of the model’s overall performance in Section 5. F1 Score and AUROC results are provided in Appendix.
Prompt Adaptation to Specific Tasks. We observe that certain prompt methods, such as GPPT, are only applicable to node classification, while GPF and GPF-plus are only suitable for graph classification. To this end, we design graph task tokens to replace the original node task tokens to make GPPT also applicable in graph-level tasks. Inspired by All-in-one and Gprompt, we utilize induced graphs in our benchmark to adapt GPF and GPF-plus for the node task. The concrete adaptation designs are shown in Appendix.
Hyperparameter Optimization. To manage the influence of hyperparameter selection and maintain fairness, the evaluation process is standardized both with and without hyperparameter adjustments. Initially, default hyperparameters as specified in the original publications are used. To ensure fairness during hyperparameter tuning, random search is then employed to optimize the settings. In each trial across different datasets, a set of hyperparameters is randomly selected from a predefined search space for each model. For additional details on metrics, default hyperparameters, search spaces, and other implementation aspects, please see Appendix.
@@ 5.1 Overall Performance of Graph Prompt Learning
In Table 2 and Table 3, we present the optimal experimental results of the supervised methods, classical ‘Pre-train & Fine-tune’ methods, and various graph prompt methods on 1-shot node/graph classification tasks across 15 datasets, where the evaluation metric is accuracy (Complete results can be referred to Appendix). Here the ’optimal’ results mean that the best result is shown from six classical ‘Pre-train & Fine-tune’ methods or from six pretraining-based variants of each graph prompt method. This presentation style can help reflect the upper bounds of classical ‘Pre-train & Fine-tune’ methods and graph prompt methods, better highlighting the advantages of graph prompts.
@@
Similar observations can also be found in 3-shot and 5-shot results in the optimal and complete versions, which can be checked in Appendix. Additionally, for comprehensive, we also offer the curve of the shot number from 1 to 10 and the accuracy in Appendix.
@@
(a) GPF-plus (1-shot node classification Task).
(b) All-in-one (1-shot graph classification Task).
@@ 5.2 Impact of Pre-training Methods
Figure 3(a) and Figure 3(b) respectively show the heat maps of Table 4 and Table 5, respectively, for help better understanding the impact of pre-training methods (Other heat maps of other graph prompt learning methods can be referred to Appendix).
@@ 5.3 Flexibility Analysis
We leverage error bound, defined by Equation 1, to measure the flexibility of graph prompt methods [22]. The error bound quantifies the difference in graph representations between the original graph and the manually manipulated graph restored by graph prompt methods. These manipulations include dropping nodes, dropping edges, and masking features. Table 5.3 presents the specific results. The experimental results indicate that all graph prompt methods significantly reduced the error
@@
compared to the original error across all types of manipulations. Notably, All-in-one and Gprompt achieved the lowest error bounds, demonstrating their flexibility in adapting to graph transformations. Specifically, All-in-one achieved a REDof 95.06, indicating a 95.06% reduction in original error, underscoring its comprehensive approach to capturing graph transformations. These findings highlight the importance of flexibility in prompt design for effective knowledge transfer. Flexible prompts that can adapt to various graph transformations enable more powerful graph data operation capability, thereby improving the performance of downstream tasks. Moreover, this flexibility is inspiring for more customized graph data augmentation, indicating more potential for other applications.
@@ 5.4 Efficiency Analysis
Besides the main observations mentioned before, we here further discuss the adaptability of graph prompts on various kinds of graphs. Specifically, considering graph prompts performance on various domains as shown in Table 3, we can find that the adaptability to various domains of graph prompt methods, particularly those of the Prompt as Token type, is not always good, where graph prompt methods even may encounter negative transfer issues. GPPT, in particular, faces significant challenges, likely because first-order subgraphs fail to comprehensively capture graph patterns, introducing noise and affecting transferability. For Large-scale Datasets like node-level dataset ogbn-arxiv in Table 2, the transferability of graph prompt methods involving subgraph operations, such as GPPT and All-in-one, requires further enhancement. Furthermore, regarding the relatively large graph-level dataset DD in Table 3, all graph prompt methods achieve positive transfer and outperform Pre-train & Fine-tune. Despite this success, the magnitude of improvement indicates that the transferability on graph-level large-scale datasets also requires further enhancement. For Homophilic & Heterophilic Datasets in Table 2, most graph prompt methods can handle both homophilic and heterophilic graph datasets effectively, with the exception of GPPT, possibly due to instability caused by the clustering algorithms of GPPT. For Pure Structural Dataset like IMDB-BINARY and COLLAB in Table 3, the ability to handle pure structural datasets of graph prompt methods needs further refinement. In detail, some ‘Pre-train & Fine-tune’ methods can achieve positive transfer on these datasets, while some variants of the All-in-one method even exhibit negative transfer. Further detailed analysis can be referred to Appendix.

### v2

@@ 3 Benchmark Methods
In general, graph prompt learning aims to transfer pre-trained knowledge with the expectation of achieving better positive transfer effects in downstream tasks compared to classical ‘Pre-train & Fine-tune’ methods. To conduct a comprehensive comparison, we first consider using a supervised method as the baseline for judging positive transfer (For example, if a ‘pre-training and fine-tuning’ approach can not beat supervised results, one negative transfer case is observed). We introduce various pre-training methods to generate pre-trained knowledge, upon which we perform fine-tuning as the comparison for graph prompts. Finally, we apply current popular graph prompt learning methods with various pre-training methods to investigate their knowledge transferability. The details of these baselines are as follows (Further introduction can be referred to Appendix A):
  • Supervised Method: In this paper, we utilize Graph Convolutional Network (GCN) [20] as the baseline, which is one of the classical and effective graph models based on graph convolutional operations [45, 61] and also the backbone for both ‘Pre-train & Fine-tune’ and graph prompt learning methods. We also explore more kinds of graph models like GraphSAGE [12], GAT [48], and Graph Transformer [38], and put the results in Appendix E.
@@ 4.1 Datasets
MUTAG 188 17.9 19.8 7 2 G small molecule
IMDB-BINARY 1000 19.8 96.53 0 2 G social network
COLLAB 5000 74.5 2457.8 0 3 G social network
@@
ENZYMES 600 32.6 62.1 3 6 G proteins
DD 1,178 284.1 715.7 89 2 G proteins
@@
To figure out how adaptive existing graph prompts on different graphs, we conduct our experiments across 15 datasets on node-level or graph-level tasks, providing a broad and rigorous testing ground for algorithmic comparisons. As shown in Table 4.1, we include 7 node classification datasets, covering homophilic datasets (Cora, Citeseer, PubMed, ogbn-arxiv) [37, 30, 14], heterophilic datasets (Wisconsin, Texas, Actor) [32, 47], and large-scale datasets (ogbn-arxiv). Additionally, we consider 8 graph classification datasets from different domains, including social networks (IMDB-B, COLLAB) [57], biological datasets (ENZYMES, PROTEINS, DD) [6, 2, 52], and small molecule datasets (MUTAG, COX2, BZR) [21, 36]. More details on these datasets can be found in Appendix C.
@@ 4.2 Implementation Details
Metrics. Following existing multi-class classification benchmarks, we select Accuracy, Macro F1 Score, and AUROC (Area Under the Receiver Operating Characteristic Curve) as our main performance metrics. The Accuracy results are used to provide a straightforward indication of the model’s overall performance in Section 5. F1 Score and AUROC results are provided in Appendix D.
Prompt Adaptation to Specific Tasks. We observe that certain prompt methods, such as GPPT, are only applicable to node classification, while GPF and GPF-plus are only suitable for graph classification. To this end, we design graph task tokens to replace the original node task tokens to make GPPT also applicable in graph-level tasks. Inspired by All-in-one and Gprompt, we utilize induced graphs in our benchmark to adapt GPF and GPF-plus for the node task. The concrete adaptation designs are shown in Appendix A.
Hyperparameter Optimization. To manage the influence of hyperparameter selection and maintain fairness, the evaluation process is standardized both with and without hyperparameter adjustments. Initially, default hyperparameters as specified in the original publications are used. To ensure fairness during hyperparameter tuning, random search is then employed to optimize the settings. In each trial across different datasets, a set of hyperparameters is randomly selected from a predefined search space for each model. For additional details on metrics, default hyperparameters, search spaces, and other implementation aspects, please see Appendix D.
@@ 5.1 Overall Performance of Graph Prompt Learning
In Table 2 and Table 3, we present the optimal experimental results of the supervised methods, classical ‘Pre-train & Fine-tune’ methods, and various graph prompt methods on 1-shot node/graph classification tasks across 15 datasets, where the evaluation metric is accuracy (Complete results can be referred to Appendix E). Here the ’optimal’ results mean that the best result is shown from six classical ‘Pre-train & Fine-tune’ methods or from six pretraining-based variants of each graph prompt method. This presentation style can help reflect the upper bounds of classical ‘Pre-train & Fine-tune’ methods and graph prompt methods, better highlighting the advantages of graph prompts.
@@
Similar observations can also be found in 3-shot and 5-shot results in the optimal and complete versions, which can be checked in Appendix E. Additionally, for comprehensive, we also offer the curve of the shot number from 1 to 10 and the accuracy in Appendix E.
@@ 5.2 Impact of Pre-training Methods
(a) GPF-plus (1-shot node classification Task).
(b) All-in-one (1-shot graph classification Task).
Figure 3(a) and Figure 3(b) respectively show the heat maps of Table 4 and Table 5, respectively, for help better understanding the impact of pre-training methods (Other heat maps of other graph prompt learning methods can be referred to Appendix E).
@@ 5.3 Flexibility Analysis
We leverage error bound, defined by Equation 1, to measure the flexibility of graph prompt methods [42]. The error bound quantifies the difference in graph representations between the original graph and the manually manipulated graph restored by graph prompt methods. These manipulations include dropping nodes, dropping edges, and masking features. Table 5.3 presents the specific results. The experimental results indicate that all graph prompt methods significantly reduced the error compared to the original error across all types of manipulations. Notably, All-in-one and Gprompt achieved the lowest error bounds, demonstrating their flexibility in adapting to graph transformations. Specifically, All-in-one achieved a REDof 95.06, indicating a 95.06% reduction in original error, underscoring its comprehensive approach to capturing graph transformations. These findings highlight the importance of flexibility in prompt design for effective knowledge transfer. Flexible prompts that can adapt to various graph transformations enable more powerful graph data operation capability, thereby improving the performance of downstream tasks. Moreover, this flexibility is inspiring for more customized graph data augmentation, indicating more potential for other applications.
@@ 5.4 Efficiency Analysis
Besides the main observations mentioned before, we here further discuss the adaptability of graph prompts on various kinds of graphs. Specifically, considering graph prompts performance on various domains as shown in Table 3, we can find that the adaptability to various domains of graph prompt methods, particularly those of the Prompt as Token type, is not always good, where graph prompt methods even may encounter negative transfer issues. GPPT, in particular, faces significant challenges, likely because first-order subgraphs fail to comprehensively capture graph patterns, introducing noise and affecting transferability. For Large-scale Datasets like node-level dataset ogbn-arxiv in Table 2, the transferability of graph prompt methods involving subgraph operations, such as GPPT and All-in-one, requires further enhancement. Furthermore, regarding the relatively large graph-level dataset DD in Table 3, all graph prompt methods achieve positive transfer and outperform Pre-train & Fine-tune. Despite this success, the magnitude of improvement indicates that the transferability on graph-level large-scale datasets also requires further enhancement. For Homophilic & Heterophilic Datasets in Table 2, most graph prompt methods can handle both homophilic and heterophilic graph datasets effectively, with the exception of GPPT, possibly due to instability caused by the clustering algorithms of GPPT. For Pure Structural Dataset like IMDB-BINARY and COLLAB in Table 3, the ability to handle pure structural datasets of graph prompt methods needs further refinement. In detail, some ‘Pre-train & Fine-tune’ methods can achieve positive transfer on these datasets, while some variants of the All-in-one method even exhibit negative transfer. Further detailed analysis can be referred to Appendix E.